# Reasoning Under Uncertainty: Exploring Probabilistic Reasoning Capabilities of LLMs

### v1

@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive performance across a wide range of Natural Language Processing (NLP) tasks, including question answering, summarization, and language understanding (Devlin et al., 2019; Raffel et al., 2020; Chowdhery et al., 2023; Joshi et al., 2017). Despite these advances, LLMs continue to exhibit notable limitations in probabilistic reasoning, as highlighted by several studies (Freedman and Toni, 2025; Gu et al., 2024; Ball et al., 2024). While there is growing interest in exploring their capabilities in probabilistic inference, it remains unclear how well these models can reason over explicit probability distributions. Existing research has primarily focused on logical reasoning, numerical problem-solving, and Bayesian inference (Chen et al., 2022; Ozturkler et al., 2022; Nafar et al., 2023, 2025; Qiu et al., 2025; Paruchuri et al., 2024), with limited attention given to reasoning over discrete probability distributions (Gu et al., 2024).
@@
To address this gap, we present the first comprehensive evaluation of the reasoning capabilities of LLMs over discrete probability distributions. We develop a structured framework that prompts models with observations drawn from a probability distribution and evaluates their responses to queries targeting either the joint or conditional distributions. Our evaluation suite consists of three complementary tasks, mode identification, maximum likelihood estimation, and sample generation, that collectively probe distinct yet interconnected aspects of probabilistic reasoning. These tasks enable us to assess whether models can perceive the core characteristics of a distribution, infer underlying parameters that govern a distribution, and generate samples that reflect target distributions. By covering both joint and conditional settings, this study captures a broad range of reasoning abilities without requiring finetuning on probabilistic datasets.
Our experiments reveal clear distinctions in the capabilities of different models, with larger models and those distilled from them significantly outperforming the smaller ones in most settings. Across all three tasks, models consistently perform better on queries about joint distributions than on conditional ones, which require deeper understanding and reasoning. Surprisingly, some models demonstrate strong sampling capabilities, generating outputs that closely align with expected distributions. We also observe that model performance varies significantly depending on how outcomes of the distributions are represented in the prompts, highlighting pertinent biases that potentially arise due to the structure of the training data utilized. In addition, models exhibit significant weaknesses in counting, particularly as the context length increases, revealing a fundamental limitation in their ability to retain and reason over extended input sequences.
To summarize, we make the following contributions in this work:
@@
  • We are the first to present an extensive evaluation of LLMs’ abilities to reason over discrete probability distributions across a variety of settings.
@@
  • We introduce a diverse suite of complementary tasks— namely, mode identification, maximum likelihood estimation, and sampling—that investigate distinct aspects of probabilistic reasoning.
  • We conduct a systematic analysis across broad-ranging LLMs, revealing a substantial performance gap between small and large models, and uncover noteworthy sampling capabilities among certain models.
@@
  • Furthermore, we highlight key limitations of current LLMs, including sensitivity to the notation of probabilistic outcomes, difficulty with accurate counting, and performance degradation when reasoning over long-context inputs.
@@ 2 Related Work
In-Context Learning: After the discovery of in-context learning (ICL) capabilities in large language models (LLMs), there has been a surge of interest in enhancing inference in zero-shot and few-shot settings. This paradigm enables models to perform tasks by conditioning on input examples without explicit parameter updates. Foundational studies have explored the mechanisms and effectiveness of ICL in various contexts (Brown et al., 2020; Min et al., 2022; Xie et al., 2021; Dai et al., 2022; Kojima et al., 2022; Wang et al., 2022). Building upon ICL, researchers have investigated methods to enable more advanced reasoning in LLMs, such as Chain-of-Thought (CoT) and Tree-of-Thoughts (ToT) frameworks that enhance problem-solving abilities (Yao et al., 2023; Wei et al., 2022).
Probabilistic Reasoning: Despite this advances, probabilistic reasoning and handling uncertainty remain challenging for LLMs (Nafar et al., 2023; Kadavath et al., 2022). Recent evaluations (Freedman and Toni, 2025) found that current models frequently violate basic probability rules such as complementarity and monotonicity, while (Ozturkler et al., 2022) proposed a structured approach, ThinkSum, decomposing probabilistic reasoning into distinct retrieval and aggregation stages. However, studies of (Gu et al., 2024) show that while LLMs understand probability concepts, they struggle to generate samples aligning with specified distributions. Another promising direction is Bayesian reasoning, where (Nafar et al., 2025) introduced the BLInD dataset to evaluate LLMs’ ability to perform Bayesian inference over interconnected variables. Furthermore, (Qiu et al., 2025) showed that training LLMs on predictions of an optimal Bayesian model significantly improves models’ capabilities.
Numerical Reasoning: Quantitative reasoning of LLMs is another aspect that has attracted notable attention. (Geva et al., 2020; Lewkowycz et al., 2022) demonstrated that numerical reasoning skills could be effectively introduced into LLMs through automated data generation and targeted training. (Chen et al., 2022) introduced the Program of Thoughts (PoT), where models express numerical reasoning as executable programs. Additionally, while (Paruchuri et al., 2024) showed that large models can reason about distributions when guided properly, (Liu et al., 2024) indicated their limited ability to infer probabilities from raw data or tables without special training and (Wang et al., 2025) highlighted biased and unreliable probability estimates by LLMs. Lastly, (Fu et al., 2024; Ball et al., 2024) revealed that LLMs struggle with counting tasks, and (Zhang et al., 2024) attributed these counting difficulties to architectural constraints and tokenization issues inherent to transformer models (Vaswani et al., 2017).
In summary, despite notable progress, key challenges remain in probabilistic reasoning, numerical accuracy, Bayesian inference, and counting, hence, addressing these gaps is essential for advancing the robustness and reliability of future LLMs.
@@ 3 Probabilistic Reasoning Framework
In this section, we introduce our framework for evaluating the probabilistic reasoning capabilities of LLMs. It begins with mathematical preliminaries, followed by details of the standardized prompt format we utilized. We then formalize three structured tasks, and describe our comprehensive evaluation pipeline. Together, these components enable a thorough and systematic comparison across different language models.
@@ 3.3 Task Definition and Performance Metrics
In this section, we describe our evaluation suite and the corresponding metrics used to assess the capability of LLMs in understanding categorical probability distributions. For each of the following tasks, we consider two settings: one involving queries about joint distributions, and the other involving queries about conditional distributions.
@@ 4 Robustness Analysis
In this section, we investigate pertinent factors that influence LLMs’ performance on probabilistic reasoning tasks by focusing on three key aspects: (1) how performance changes as the task difficulty increases, (2) how sensitive models are to variations in the label notation used to represent outcomes of the random variables, and (3) how model behavior differs when input prompts include direct, raw sample observations instead of empirical frequencies. These aspects expose distinct challenges for models and offer deeper insights into their strengths and weaknesses.
@@ 5 Experiments
This section outlines the experimental setup used to evaluate LLMs on probabilistic reasoning tasks. It introduces the models selected for evaluation and describes key implementation details. Results are presented for the three core tasks, followed by additional studies that examine factors influencing model performance.
@@ 5.1 Model Selection
In this study, we evaluate several instruction-finetuned LLMs, selected to represent a wide range of model sizes and capabilities. The models evaluated include Llama3.1-8B, Llama3.3-70B (Grattafiori et al., 2024), Qwen2.5-7B (Yang et al., 2024), DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), GPT-4o-mini (Hurst et al., 2024), and OpenAI’s recently published model, GPT-4.1-mini (Achiam et al., 2023). These models span a broad spectrum of parameter sizes, from 7B to 70B, including specialized reasoning models, offering a comprehensive comparison of how different architectures handle various settings.
@@ 5.2 Experimental Setup
In our experiments, we follow the standard naming convention for random variables, using names such as $X,Y,Z,T,$ etc., and set $K:=5\times|\mathcal{L}|$, meaning the number of observations is five times larger than the joint space. We then select $|\mathcal{L}|$ values that sum to $K$, as frequencies, and randomly assign them to different joint outcomes. However, as discussed in Section  4.2, models are sensitive to how these frequencies are linked to specific outcomes, even when both the set of labels and frequency values are unchanged. To reduce the impact of this sensitivity, we create ten different ways of assigning frequencies to joint outcomes and generate ten prompts for each assignment, resulting in 100 prompts per task. This helps ensure that variations in outcome-frequency mapping minimally impacts the evaluation results. For larger $|\mathcal{L}|$, we carefully select the frequency values to generate a distribution with entropy comparable to that of the simpler distributions, ensuring consistency across experiments. Given the models’ constraints in handling raw samples in the input, we incorporate the empirical frequency of observed samples in most experiments and analyze the comparison between the two approaches in Section 5.4.2.
@@ 5.3 Task-Specific Experimental Results
This section presents a detailed analysis of the experimental results for each task. To evaluate performance on conditional queries, one random variable $X_{c}$ and one of its possible outcomes $x_{c}$ are randomly selected as evidence. Another variable is then chosen as the query variable $X_{q}$ and the model is asked to solve one of the tasks based on the conditional distribution $P(X_{q}\mid X_{c}=x_{c})$. For all tasks, each model is prompted 100 times using 10 different prompts, and the outputs are parsed to extract the final answer for each query.
@@
Mode Identification: In this task, models are asked to identify the mode of either the joint distribution (joint-mode) or a conditional distribution (cond-mode). The joint-mode task requires selecting the outcome with the highest overall frequency, without additional computation. In contrast, the cond-mode task involves identifying the most frequent outcome given a specified condition, requiring models to filter and reason over a subset of the distribution. Table 1 reports model accuracy as the joint space size $|\mathcal{L}|$ increases. Although the cond-mode task involves fewer possible outputs, the added conditional reasoning makes it more challenging. Consequently, most models achieve slightly lower accuracy on cond-mode compared to joint-mode. While smaller models struggle as the distribution grows, Llama3.3-70B and GPT4.1-mini maintain near-perfect accuracy across all joint sizes. The details of different distributions can be found in Appendix B.
@@
Maximum Likelihood Estimation: In this task, models are asked to predict either the probability of all joint outcomes (joint-MLE) or the conditional probabilities $P(X_{q}=x_{q}\mid X_{c}=x_{c})$ for each $x_{q}\in\mathcal{L}_{q}$ (cond-MLE). For evaluation, the TVD is computed for each prompt, and the average is reported over 100 prompts. As shown in Table 2, the joint-MLE task appears to be straightforward for all models, with even smaller models achieving near-zero TVD across all joint sizes. In contrast, the cond-MLE task is more challenging, as it requires models to derive conditional probabilities from the joint distribution. This added complexity results in consistently higher TVD values, with a clear performance gap between smaller and larger models.
Sample Generation: This task evaluates the generative abilities of LLMs by requiring them to generate $K^{\prime}$ samples autoregressively from either the joint distribution (joint-sampling) or a conditional distribution (cond-sampling). We set $K^{\prime}:=7\times|\mathcal{L}|$ to ensure that the generated samples are sufficient to approximate the target distribution. Since it is inherently difficult to reproduce a target empirical distribution with a finite number of samples, we also compute the TVD between samples drawn from Python’s random library and the true distribution. This serves as a baseline for comparing the quality of samples generated by the models.
@@
As shown in Figure 2, TVD in the joint-sampling task increases with $|\mathcal{L}|$. Interestingly, Llama3.3-70B and GPT4.1-mini achieve TVD values even lower than those from Python’s random sampler, which includes sampling noise, suggesting that the underlying distribution of their generated samples is remarkably close to the target. In contrast, all models struggle with the cond-sampling task, as it implicitly requires the models to first estimate the conditional probabilities and then generate samples that align with the resulting distribution. Surprisingly, even though no model reaches a satisfactory level of performance, DeepSeek-R1-Distill-Qwen, designed for reasoning tasks, achieves the best performance on this task, outperforming even the strongest models in other settings.
Beyond distributional evaluation, we also assessed whether model-generated samples exhibit true independence. Using the Durbin–Watson statistic (Durbin and Watson, 1951) and transition matrix entropy (Shannon, 1948), we found evidence of correlations in all models’ outputs, indicating that generations are not fully independent and are influenced by previously produced samples, consistent with the autoregressive nature of LLMs. In a related experiment, GPT-4.1-mini was prompted to generate one sample at a time across multiple prompts, we observed that it predominantly produced high-frequency outcomes while almost entirely avoiding rare ones. These findings highlight that current LLMs not only struggle with conditional sampling but also face challenges in generating truly independent samples. Full experimental details and results are provided in Appendix D.
To validate our findings on real-world data, we evaluated all six task configurations using the Mushroom dataset (mus, 1981). We focused on a subset of this dataset, resulting in a highly skewed distribution. The detailed results for each task are provided in Appendix E, enabling a comprehensive comparison of model performance under realistic conditions.
@@ 5.4 Robustness Analysis Results
This section presents the results of analyzing two key aspects: the sensitivity of models to variations in outcomes of the distributions, and the effect of providing raw samples in the prompt instead of their empirical frequencies.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive performance across a wide range of Natural Language Processing (NLP) tasks, including question answering, summarization, and language understanding [7, 30, 5, 17]. Despite these advances, LLMs continue to show notable limitations in probabilistic reasoning, as highlighted by several studies [9, 13, 2]. Existing research has largely examined LLMs’ abilities in logical reasoning, numerical problem-solving, and Bayesian inference [4, 27, 24, 25, 29], but their ability to understand probability distributions has received far less attention [13, 28]. Exploring this capability is crucial, as reasoning over distributions lies at the core of many tasks involving uncertainty, prediction, and decision-making [16, 21, 31]. This gap leads to an important question: when provided with samples drawn from a distribution, can an LLM identify its structure, estimate its parameters, and produce responses consistent with that distribution?
@@
To address this, we develop a structured framework that prompts models with observations drawn from the joint probability distribution of discrete random variables and evaluates their responses to queries about joint and conditional distributions. Our evaluation suite consists of three complementary tasks, mode identification, maximum likelihood estimation, and sample generation. Together, these tasks probe distinct but interconnected skills: recognizing the most likely outcomes, estimating underlying probabilities, and producing samples that align with the target distribution. By covering both joint and conditional settings, our study provides a broad view of how well LLMs can approximate key features of probability distributions without requiring finetuning on probabilistic datasets.
Our experiments reveal several important trends. Larger models, and those distilled from them, significantly outperform smaller ones in most settings. Across all three tasks, models perform better on joint distributions than on conditional ones, which demand deeper reasoning and additional computation. Some models show surprising strengths in sampling, generating outcomes that closely match the expected distribution. On the other hand, we identify persistent weaknesses. Performance is highly sensitive to how outcomes of the distributions are represented in the prompt, suggesting that even superficial variations in notation can significantly influence reasoning. Counting also remains a fundamental challenge: as the number of samples in the context grows, models lose track of frequencies, leading to inaccurate estimates. Conditional reasoning is particularly challenging in zero-shot settings, but providing a simple in-context example improves the performance significantly. More broadly, while LLMs can often handle probabilistic tasks using their background knowledge, they struggle as the tasks become more complex. In these cases, they need additional support, whether through external tools like a code interpreter or in-context examples, to achieve reliable results.
In summary, this work makes the following contributions:
@@
  • We present the first large-scale evaluation of LLMs’ ability to understand discrete probability distributions across a variety of settings.
@@
  • We introduce a suite of complementary tasks—mode identification, maximum likelihood estimation, and sampling—that capture distinct aspects of probabilistic reasoning.
  • We systematically compare a broad range of LLMs, revealing substantial performance gaps across model sizes and unexpected strengths in sampling.
@@
  • We highlight key limitations, including sensitivity to notation of probabilistic outcomes, difficulties with counting, and challenges with long-context reasoning.
@@ 2 Related Work
In-Context Learning: After the discovery of in-context learning (ICL) capabilities in large language models (LLMs), there has been a surge of interest in enhancing inference in zero-shot and few-shot settings. This paradigm enables models to perform tasks by conditioning on input examples without explicit parameter updates. Foundational studies have explored the mechanisms and effectiveness of ICL in various contexts [3, 23, 37, 6, 19, 35]. Building upon ICL, follow-up work has introduced methods such as Chain-of-Thought (CoT) and Tree-of-Thoughts (ToT) [39, 36], which aim to extend LLMs’ reasoning capabilities and improve performance on more complex problems.
Probabilistic Reasoning: Despite these advances, probabilistic reasoning and handling uncertainty remain challenging for LLMs [24, 18]. Recent evaluations [9] found that current models frequently violate basic probability rules such as complementarity and monotonicity, and [27] proposed a structured approach, ThinkSum, decomposing probabilistic reasoning into distinct retrieval and aggregation stages. Other studies show that while LLMs understand probability concepts, they struggle to generate samples aligning with specified distributions [13]. Bayesian reasoning has also emerged as a promising direction. The BLInD dataset was introduced to evaluate LLMs’ ability to perform Bayesian inference [25], and training LLMs to imitate an optimal Bayesian model has been shown to improve their probabilistic reasoning capabilities [29].
Numerical Reasoning: Another important area of study is LLMs’ quantitative reasoning. Several works have shown that numerical reasoning skills can be effectively enhanced through automated data generation and targeted training [11, 20]. The Program of Thoughts (PoT) approach expresses numerical reasoning as executable programs, enabling models to solve complex calculations [4]. While some studies have shown that large models can reason about continuous distributions when properly guided [28], others highlight that LLMs struggle to infer probabilities from raw data or tables without special training [22], and their probability estimates can be biased or unreliable [34]. Counting tasks also proved to be challenging [10, 2], with difficulties partly attributed to architectural constraints and tokenization issues inherent to transformers [40, 33].
In summary, while LLMs have shown promising capabilities in probabilistic reasoning and Bayesian inference, their ability to understand discrete probability distributions, especially joint and conditional relationships, remains largely unexplored. This gap motivates our work, where we provide a focused and extensive study of LLMs’ understanding of discrete joint and conditional probability distributions.
@@ 3 Probabilistic Reasoning Framework
In this section, we present our framework for evaluating LLMs’ capability to understand probability distributions. We begin with the mathematical preliminaries and describe the prompt format used in our experiments. Then, we formalize the three structured tasks that form the core of our evaluation, and finally present our comprehensive evaluation pipeline. In this work, we decided to focus on discrete probability distributions, as methods and evaluation strategies for continuous distributions differ enough that addressing both types in one paper would risk superficial coverage.
@@ 3.3 Task Definition and Performance Metrics
In this section, we describe our three structured tasks and the metrics used to assess LLMs’ performance on them. We consider two settings: one focused on queries about joint distributions and the other on queries about conditional distributions.
@@ 4 Robustness Analysis
In this section, we examine pertinent factors that influence LLMs’ performance on probabilistic reasoning tasks, focusing on three key aspects: (1) how performance changes with increasing task difficulty, (2) how sensitive models are to variations in the notation used to represent random variable outcomes, and (3) how models behavior differs when prompts provide raw sample observations rather than empirical frequencies. Together, these aspects highlight distinct challenges for LLMs and provide deeper insight into their strengths and limitations.
@@ 5 Experiments
This section presents the experimental setup for evaluating LLMs on probabilistic reasoning tasks. We introduce the models evaluated, describe key implementation details, and report results for the three core tasks. We then present additional analyses that explore factors influencing model performance.
@@ 5.1 Model Selection
In this study, we evaluate several instruction-finetuned LLMs, selected to represent a wide range of model sizes and capabilities. The models evaluated include Llama3.1-8B, Llama3.3-70B [12], Qwen2.5-7B [38], DeepSeek-R1-Distill-Qwen-7B [14], GPT-4o-mini [15], and GPT-4.1-mini [26]. These models span a broad spectrum of parameter sizes, from 7B to 70B, offering a comprehensive comparison of how different architectures handle various settings.
@@ 5.2 Experimental Setup
In our experiments, we follow the standard naming convention for random variables, using names such as $X,Y,Z$ and use English letters for the outcomes of the random variables . We set $K:=5\times|\mathcal{L}|$, meaning the number of observations is five times larger than the joint space. We then select $|\mathcal{L}|$ values that sum to $K$, as frequencies, and randomly assign them to different joint outcomes. However, as discussed in Section  4.2, models are sensitive to how these frequencies are linked to specific outcomes, even when both the set of labels and frequency values are unchanged. To reduce the impact of this sensitivity, we create ten different ways of assigning frequencies to joint outcomes and generate ten prompts for each assignment, resulting in 100 prompts per task. This helps ensure that variations in outcome-frequency mapping minimally impacts the evaluation results. For larger $|\mathcal{L}|$, we carefully select the frequency values to generate a distribution with entropy comparable to that of the simpler distributions, ensuring consistency across experiments. Given the models’ constraints in handling raw samples in the input, we incorporate the empirical frequency of observed samples in most experiments and analyze the comparison between the two approaches in Section 5.4.2.
@@ 5.3 Task-Specific Experimental Results
This section presents a detailed analysis of the experimental results for each task. To evaluate performance on conditional queries, one random variable $X_{c}$ and one of its possible outcomes $x_{c}$ are randomly selected as evidence. Another variable is then chosen as the query variable $X_{q}$ and the model is asked to solve one of the tasks based on the conditional distribution $P(X_{q}\mid X_{c}=x_{c})$. For all tasks, each model is prompted 100 times using 10 different prompts, and the outputs are parsed to extract the final answer for each query. The details of different distributions can be found in Appendix B.
@@
Mode Identification: In this task, models are asked to identify the mode of either the joint distribution (joint-mode) or a conditional distribution (cond-mode). The joint-mode task requires selecting the outcome with the highest overall frequency, without additional computation. In contrast, the cond-mode task involves identifying the most frequent outcome given a specified condition, requiring models to filter and reason over a subset of the distribution. Table 1 reports model accuracy as the joint space size $|\mathcal{L}|$ increases. Although cond-mode involves fewer possible outputs, its added conditional reasoning makes it more challenging, leading to slightly lower accuracy than joint-mode. While Llama3.3-70B and GPT4.1-mini remain near-perfect, smaller models degrade as the joint size grows. Overall, although models can handle simpler instances of these tasks using their pretraining knowledge, their accuracy drops noticeably as tasks become more complicated. To mitigate this, we conducted additional experiments using one-shot prompting instead of the default zero-shot setting. The results show a substantial improvement in performance, underscoring the effectiveness of few-shot learning for these tasks. The details of these experiments can be found in Appendix D.
@@
Maximum Likelihood Estimation: In this task, models are asked to predict either the probability of all joint outcomes (joint-MLE) or the conditional probabilities $P(X_{q}=x_{q}\mid X_{c}=x_{c})$ for each $x_{q}\in\mathcal{L}_{q}$ (cond-MLE). For evaluation, the TVD is computed for each prompt, and the average is reported over 100 prompts. As shown in Table 2, the joint-MLE task is relatively straightforward for all models, with even smaller models achieving acceptable TVD values across all joint sizes. In contrast, the cond-MLE task is more challenging, as it requires models to derive conditional probabilities from the joint distribution, leading to consistently higher TVD values and a clear gap between smaller and larger models. To mitigate this challenge, we also ran one-shot prompting experiments, which significantly reduced TVD values, showing that even a simple example can boost performance on probabilistic reasoning tasks. Full details of these experiments are provided in Appendix D.
@@
Sample Generation: This task evaluates the generative abilities of LLMs by requiring them to generate $K^{\prime}$ samples from either the joint distribution (joint-sampling) or a conditional distribution (cond-sampling). We set $K^{\prime}:=7\times|\mathcal{L}|$ to ensure that the generated samples are sufficient to approximate the target distribution. Since it is inherently difficult to reproduce a target empirical distribution with a finite number of samples, we also compute the TVD between $K^{\prime}$ samples drawn from Python’s random library and the true distribution and use it as a baseline to compare the quality of samples generated by the models. As shown in Figure 2, TVD in the joint-sampling task increases with $|\mathcal{L}|$. Interestingly, Llama3.3-70B and GPT4.1-mini achieve TVD values even lower than those from Python’s random sampler, which includes sampling noise, suggesting that the underlying distribution of their generated samples is remarkably close to the target. In contrast, all models struggle with the cond-sampling task, as it implicitly requires the models to first estimate the conditional probabilities and then generate samples that align with the resulting distribution. Surprisingly, even though no model reaches an acceptable performance, DeepSeek-R1-Distill-Qwen, designed for reasoning tasks, outperforms other strong models.
Beyond distributional evaluation, we also examined whether model-generated samples were truly independent. Using the Durbin–Watson statistic and transition matrix entropy [8, 32], we found correlations in all models’ outputs. In a related experiment, when GPT-4.1-mini was prompted to generate one sample per prompt, it strongly favored high-frequency outcomes while rarely producing rare ones. These results suggest that while capable LLMs can generate samples that closely match target empirical probabilities, they struggle to produce independent samples, consistent with the autoregressive nature of LLMs. To further validate our findings, we repeated all six task configurations on a highly-skewed subset of the Mushroom dataset [1], with results aligning closely to those from synthetic data (see Appendices E,  F).
@@ 5.4 Robustness Analysis Results
This section presents the analysis of two key aspects: the sensitivity of models to variations in outcomes of the random variables, and the effect of providing raw samples in the prompt instead of their empirical frequencies. These analyses shed light on how superficial changes and input formats can significantly affect performance.